OpenStudy (anonymous):

the ratio of men to woman in a certain factory is 3 to 4. how many workers are there?

OpenStudy (anonymous):

So you know there are 210 men, so you would have to do \[\frac{ 3 }{ 4 } = \frac{ 210 }{ x }\] \[x = \frac{ 4*210 }{ 3}\] So what do you think it is?

OpenStudy (anonymous):

280?

OpenStudy (anonymous):

Correct, now is that one of your choices?